Translate the following problems into algebraic equations:


1) In a certain triangle the measure of one angle is double the measure of a second angle but is 10 degrees less than the measure of the third angle. [The sum of the measures of three interior angles of a triangle is always 180 degrees.] What is the measure of each angle?
2) Tom’s uncle owns a triangular piece of land. The perimeter fence that surrounds the land measures 378 yards. The shortest side is 30 yards longer than one-half of the longest side. The second longest side is 2 yards shorter than the longest side. What is the length of each side?

Answer by stanbon(75887)

1) In a certain triangle the measure of one angle is double the measure of a second angle but is 10 degrees less than the measure of the third angle. [The sum of the measures of three interior angles of a triangle is always 180 degrees.] What is the measure of each angle?
-------------------------
Let the 2nd angle be "x"
Then "one angle" is "2x"
3rd angle is 2x+10
----------------------
EQUATION
x + 2x + 2x+10 = 180
5x+10 = 180
5x = 170
-----------------
x = 34 (2nd angle)
2x = 68 (one angle)
2x+10 = 78 (3rd angle)
------------------------
Cheers,
Stan H.





2) Tom’s uncle owns a triangular piece of land. The perimeter fence that surrounds the land measures 378 yards. The shortest side is 30 yards longer than one-half of the longest side. The second longest side is 2 yards shorter than the longest side. What is the length of each side?